# Zero-shot sampling of adversarial entities in biomedical question answering

### v1

@@ 5 Attack analysis and diagnostics
Figure 4: (Top) An entity subsitution attack using n-acetylglucosamine as the replacement entity for metoprolol creates an adversarial distractor. (Bottom) Heatmaps of token-wise Shapley values for a question before and after the adversarial attack on option C. The model prediction changes from the correct choice of B (unperturbed) to the incorrect choice of D (adversarially perturbed).
Entity substitution attacks manipulate explainability.
@@
Post-hoc analysis is a common way to understand the rationale behind model decisions (Murdoch et al., 2019). Score-based feature importance such as that based on the Shapley values are routinely used to construct post-hoc explanation of model prediction through feature attribution (Chen et al., 2023). The stability of model explanation is a growing concern for their proper use (Alvarez-Melis and Jaakkola, 2018; Hancox-Li, 2020; Lin et al., 2023). Investigations on the susceptibility of LLM explanations to adversarial perturbation in QA are still lacking. For score-based feature importance, an explainer $\phi(\cdot)$ maps the $m$-dimensional feature $\mathbf{x}\in\mathbb{R}^{m}$ to scores $\phi(\mathbf{x})\in\mathbb{R}^{m}$. In the current context, we assess the changes in feature importance represented by Shapley values (Chen et al., 2023) before ($\phi(\mathcal{Q})$) and after ($\phi(\mathcal{Q}^{\prime})$) the adversarial attack by entity substitution $\mathcal{Q}\rightarrow\mathcal{Q}^{\prime}$. For convenience of discussion, we use Shapley profile to refer to $\phi(\mathcal{Q})$ (or $\phi(\mathcal{Q}^{\prime})$), the token-wise Shapley values calculated over $\mathcal{Q}$ (or $\mathcal{Q}^{\prime}$).
@@
We investigated the changes in the Shapley profile in the two regimes of adversarial entities for open-source LLMs, because the calculation requires access to model weights. We found that the token with the largest Shapley value (top-1 feature) is highly correlated with the model’s prediction before and after adversarial perturbation, with an example shown in Fig. 4 produced using Flan-T5-large. This behavior is an indicator of the symbol-binding characteristic of LLMs in multiple-choice QA (Robinson and Wingate, 2023). Moreover, the Shapley profile typically exhibits pronounced changes before and after entity perturbation, especially in the text locations close to the perturbed entities (Fig. 4 and Appendix E). Together, these observations indicate that the substitution attacks effectively manipulate model explanations, because the explanation after attack aligns with the wrong prediction, as if they were correct. The behavior occurs despite unchanged context, stem, and key ($\mathcal{C}$, $\mathcal{S}$, and $k$), which should contain the most important information for understanding and answering the question (Chai and Jin, 2004).
6 Related works
Text substitution attack
Adversarial attacks in neural language models have been demonstrated from the character to the sentence level (Zhang et al., 2020). The majority of text substitution attacks employ typos (Gao et al., 2018) and synonyms (Mozes et al., 2021), which effectively impose lexical constraints. Word frequency has been used as a metric to select adversarial examples in (Mozes et al., 2021). Synonym substitution attacks have been demonstrated in sentiment analysis (Liu et al., 2022). Alzantot et al. (2018) used a genetic algorithm to search for adversarial word substitution. Model sensitivity to entity-level perturbations such as substitution and swapping create problematic situations for knowledge-intensive tasks, such as knowledge conflicts (Longpre et al., 2021), compromised performance in machine reading comprehension (Yan et al., 2022) and table interpretation tasks (Koleva et al., 2023).
Robustness in QA
The robustness of QA systems has been studied using perturbative methods (Jia and Liang, 2017) and meaning-preserving transformations (Gan and Ng, 2019; Elazar et al., 2021). Sen and Saffari (2020) introduced a set of perturbations to assess language model generalization between QA datasets. Richardson and Sabharwal (2020) proposed a distractor design method using proximity in semantic space to probe model understanding of word senses. Awadalla et al. (2022) examined the connection between in- and out-of-distribution robustness for different LLMs. The performance of LLMs is sensitive to the ordering of options (Pezeshkpour and Hruschka, 2023; Zheng et al., 2023). Overall, the brittleness of knowledge in LLMs remains a primary limitation for their applications (Razniewski et al., 2021; Elazar et al., 2021; Ji et al., 2023).
Adversarial attack characteristics
The defining characteristic of adversarial attacks is to induce false predictions in trained models. Besides, attacks may possess transferability, universality (Zou et al., 2023), and they may obstruct model explanation (Noppel and Wressnegger, 2024). Transferability refers to the effectiveness of the same example being adversarial to a number of distinct models (Papernot et al., 2016). Universal adversarial perturbations are reusable across many examples tested on the same model (Moosavi-Dezfooli et al., 2017; Wallace et al., 2019a). Model explanations may be manipulated by specifically designed perturbations to the inputs (Dombrowski et al., 2019) or biased data subsampling (Laberge et al., 2023).
@@ 7 Discussion and future work
The present study illustrates that side information introduced through entity substitution leads to noticeable degradation of model performance, by modifying only a single sampling step and without gradient computation. Moreover, type-constrained entity substitution makes sure that the replacement entity appears natural to human and is less likely to be detected using grammar check or topical filtering (Morris et al., 2020). The setting simulates realistic scenarios where the “attacks" may be initiated by unsuspecting users, such as biomedical professionals using an LLM-based clinical decision support system (Liu et al., 2023). In the examples illustrated in Fig. 4 and Appendix E, adversarial distractors can lead to disease misdiagnosis or prescription of the wrong medication, which are detrimental scenarios that may occur in algorithmic decision-making in biomedicine. Further improvement on the attack performance may be achieved by sequential sampling, an intrinsic approach to boost the sampling performance, or by changing the entity vocabulary of the perturbation set, an extrinsic approach that depends on the user and application setting. Our approach may be integrated into interactive platforms for adversarial testing in open-ended human-machine conversation (Chao et al., 2023) for online model auditing.
Future research may focus on how to make LLMs more reliable through prompt engineering (Si et al., 2023; Nori et al., 2023) or leveraging retrieval from nonparametric knowledge bases (Weikum et al., 2021; Soman et al., 2023) or text sources (Wang et al., 2023c) to improve model generalization to long-tail entities. The DWS methodology may be extended to kernel methods (Hofmann et al., 2008) or metric learning (Bellet et al., 2015), which can capture more general distance dependence. Besides vulnerability identification, the current work motivates adversarial defense strategies based on generalized distances to establish a tiered system for validating adversarial robustness in user-facing and domain-oriented applications.
Limitations
Although NEs convey parts of the key information in the question, they fall short of capturing the entirety of the textual meaning. To explore other attack surfaces within the adversarial landscape of QA, one would need to evaluate the robustness in multidimensional perturbation space (Wang et al., 2022). Besides, the meaning of Shapley values calculated here requires additional interpretation due to feature correlation. Although we use them in part to reveal the importance of specific tokens, their robustness is also worth exploring using sentence-level features (Chen et al., 2023).
Ethics statement
The scenarios considered in this work using public datasets relate to potential real-world deployment of LLMs in biomedical education and clinical decision support. Our results show that the lack of adversarial testing methods in high-stakes domains may incur serious consequences because of naturally looking adversarial examples. The design and deployment of adversarial testing should be specific to and carried out at every stage of the model development cycle, paired with tools such as adversarially robust model explainability methods for understanding model behavior.
Appendix A Adversarial distractor generation
@@
$\mathrm{argmax}$ $\Pr(Y=\mathcal{K}|\mathcal{Q}(d_{i,j}\rightarrow d_{i,j}^{\prime}))$
$\neq\mathrm{argmax}\Pr(Y=\mathcal{K}|\mathcal{Q}).$ (8)
@@ Appendix B Proof of the sampling regimes
Here we present a simple proof of the deterministic limit of the powerscaled distance-weighted sampling described in Eq. (3) of the main text. Although the distance function $h$ is not given a specific form, we assume it to satisfy the basic properties of a metric. Given a perturbation set -$\mathcal{E}_{\textrm{-}k}=\{e_{1},...,e_{m}\}$ with respect to a key $k$, there exists a corresponding set of distance values -$\mathcal{H}_{\textrm{-}k}$.
@@ B.3 The case of $n=0$
When $n=0$, it is straightforward to show that $p_{j}|_{n=0}=1/m$, regardless of $j$ using the definition in Eq. (3). This scenario therefore corresponds to uniform random sampling.
Answer the following question without explanation. [[[[Content]]]]: A 23-year-old pregnant woman at 22 weeks gestation presents with burning upon urination. She states it started 1 day ago and has been worsening despite drinking more water and taking cranberry extract. She otherwise feels well and is followed by a doctor for her pregnancy. Her temperature is 97.7°F (36.5°C), blood pressure is 122/77 mmHg, pulse is 80/min, respirations are 19/min, and oxygen saturation is 98% on room air. Physical exam is notable for an absence of costovertebral angle tenderness and a gravid uterus. [[[[Question]]]]: Which of the following are the best treatment for this patient? A: Ampicillin B: Ceftriaxone C: Doxycycline D: Nitrofurantoin [[[[Answer]]]]:
@@ Appendix D Prompt formatting
We structured the prompt with bracketed tags to divide different parts (Context, Question, Answer). An example taken from the MedQA-USMLE dataset (Jin et al., 2021) is shown below. For the MedMCQA dataset (Pal et al., 2022), only the Context and Answer [^1] Using the Context tag improves model performance than the Question tag here. tags were used because the questions are short and generally contain no context. The prompt formatting improves model performance and facilitates answer parsing in the model-generated text. The increased space between paragraphs in the following is only to assist viewing and is not included in the actual prompt.
@@ Appendix E Model evaluation and attack examples
We list some examples of successful entity substitution attacks using adversarial distractors following Fig. 4 in the main text. For each example, the text box on top (the prompt formatting removed) shows the original options on the left side and the perturbed options on the right side. The entity of interest before and after perturbation is colored in red and blue, respectively. The checkmark indicates the model prediction on unperturbed and perturbed questions. Beneath the text box are the corresponding heatmaps of token-wise Shapley values (or Shapley profile) before and after the entity substitution attack. The Shapley profiles for the MedQA-USMLE dataset (Jin et al., 2021) were generated using the model Flan-T5-large and those for the MedMCQA (Pal et al., 2022) using the model Flan-T5-xxl (Chung et al., 2022). The Shapley values were calculated using the partition explainer in shap, which uses the Owen values as an approximation (Chen et al., 2023).
Of the LLMs evaluated in this work, the proprietary models (GPT-3.5, PaLM2) were accessed through third-party APIs, while the open-source models were accessed on local cache to improve speed. The evaluations were carried out on a variety of NVIDIA GPUs depending on the model size. For smaller model such as the Flan-T5 series (Chung et al., 2022), the runtime on each type of questions (involving drug or disease mentions) in a dataset was generally between 10 and 30 minutes, while it was from one to more than four hours for larger model such Flan-UL2 (Tay et al., 2022) and Palmyra-Med-20B (Kamble and Alshikh, 2023). For proprietary models accessed through their APIs, the runtime for each type of questions is between two and four hours.
An 18-year-old man presents with a sudden loss of consciousness while playing college football. There was no history of a concussion. Echocardiography shows left ventricular hypertrophy and increased thickness of the interventricular septum. Which is the most likely pathology underlying the present condition? A: Mutation in the myosin heavy chain   ✓ A: Mutation in the myosin heavy chain B: Drug abuse B: Drug abuse C: Viral infection C: Muller barth menger syndrome   ✓ D: Autoimmunity of myocardial fibers D: Autoimmunity of myocardial fibers
@@
A 41-year-old woman comes to the physician because of a 3-month history of anxiety, difficulty falling asleep, heat intolerance, and a 6-kg (13.2-lb) weight loss. The patient’s nephew, who is studying medicine, mentioned that her symptoms might be caused by a condition that is due to somatic activating mutations of the genes for the TSH receptor. Examination shows warm, moist skin and a 2-cm, nontender, subcutaneous mass on the anterior neck. Which of the following additional findings should most raise concern for a different underlying etiology of her symptoms? A: Nonpitting edema   ✓ A: Nonpitting edema B: Atrial fibrillation B: Atrial fibrillation C: Lid lag C: Lid lag D: Fine tremor D: Nonpuerperal galactorrhea   ✓
@@
A 25-year-old man is brought to the emergency department after his girlfriend discovered him at home in a minimally responsive state. He has a history of drinking alcohol excessively and using illicit drugs. On arrival, he does not respond to commands but withdraws all extremities to pain. His pulse is 90/min, respirations are 8/min, and blood pressure is 130/90 mm Hg. Pulse oximetry while receiving bag-valve-mask ventilation shows an oxygen saturation of 95%. Examination shows cool, dry skin, with scattered track marks on his arms and legs. The pupils are pinpoint and react sluggishly to light. His serum blood glucose level is 80 mg/dL. The most appropriate next step in management is intravenous administration of which of the following? A: Naloxone A: Naloxone B: Phentolamine   ✓ B: Phentolamine C: Methadone C: Methadone   ✓ D: Naltrexone D: Tienilic acid
@@
A 70-year-old woman presents to the office for a yearly physical. She states she has recently started experiencing pain in her legs and her back. Last year, she experienced a fracture of her left arm while trying to lift groceries. The patient states that she does not consume any dairy and does not go outside often because of the pain in her legs and back. Of note, she takes carbamazepine for seizures. On exam, her vitals are within normal limits. You suspect the patient might have osteomalacia. Testing for which of the following is the next best step to confirm your suspicion? A: 25-hydroxyvitamin D   ✓ A: 25-hydroxyvitamin D B: 1,25-hydroxyvitamin D B: 1,25-hydroxyvitamin D C: Pre-vitamin D3 C: Pre-vitamin D3   ✓ D: Dietary vitamin D2 D: Stiripentol
@@
Which of the following is granted orphan drug status for the treatment of Dravet’s syndrome? A: Stiripentol A: Stiripentol B: Icatibant B: Tocilizumab   ✓ C: Pre-vitamin D3 C: Pitolisant D: Tafamidis   ✓ D: Tafamidis
@@
Which among the following may be used as a sedatives-hypnotic? A: Zolmitriptan A: Zolmitriptan   ✓ B: Zileuton B: Zileuton C: Zolpidem   ✓ C: N-acetylglucosamine D: Zalcitabine D: Zalcitabine

### v2

@@ 5.2 Scaling characteristics from adversarial entities
Figure 5: (Top) An entity substitution attack using n-acetylglucosamine as the replacement entity for metoprolol creates an adversarial distractor. (Bottom) Heatmaps of token-wise Shapley values for a question before and after the adversarial attack on option C. The model prediction changes from the correct choice of B (unperturbed) to the incorrect choice of D (adversarially perturbed).
5.3 Adversarial entities manipulate explainability
@@
Post-hoc analysis is a common way to understand the rationale behind model decisions (Murdoch et al., 2019). Score-based feature importance such as that based on the Shapley values is routinely used to construct post-hoc explanation of model prediction through feature attribution (Chen et al., 2023). The stability of model explanation is a growing concern for their proper use (Alvarez-Melis & Jaakkola, 2018; Hancox-Li, 2020; Lin et al., 2023). Investigations on the susceptibility of LLM explanations to adversarial perturbation in QA are still lacking. For score-based feature importance, an explainer $\phi(\cdot)$ maps the $m$-dimensional feature $\mathbf{x}\in\mathbb{R}^{m}$ to scores $\phi(\mathbf{x})\in\mathbb{R}^{m}$. In the current context, we assess the changes in feature importance represented by Shapley values (Chen et al., 2023) before ($\phi(\mathcal{Q})$) and after ($\phi(\mathcal{Q}^{\prime})$) the adversarial attack by entity substitution $\mathcal{Q}\rightarrow\mathcal{Q}^{\prime}$. For convenience of discussion, we use Shapley profile to refer to $\phi(\mathcal{Q})$ (or $\phi(\mathcal{Q}^{\prime})$), the token-wise Shapley values calculated over $\mathcal{Q}$ (or $\mathcal{Q}^{\prime}$).
@@
We investigated the changes in the Shapley profile in the two regimes of adversarial entities for open-source LLMs, because the calculation requires access to model weights. We found that the token with the largest Shapley value (top-1 feature) is highly correlated with the model’s prediction before and after adversarial perturbation, with an example shown in Fig. 5 produced using Flan-T5-large. This behavior is an indicator of the symbol-binding characteristic of LLMs in multiple-choice QA (Robinson & Wingate, 2023). Moreover, the Shapley profile typically exhibits pronounced changes before and after entity perturbation, especially in the text locations close to the perturbed entities (Fig. 5 and Appendix E). Together, these observations indicate that the substitution attacks effectively manipulate model explanations, because the explanation after the attack aligns with the wrong prediction, as if they were correct. The behavior occurs despite unchanged context, stem, and key ($\mathcal{C}$, $\mathcal{S}$, and $k$), which should contain the most important information for understanding and answering the question (Chai & Jin, 2004).
@@ 6 Discussion and future work
The present study illustrates that distracting information introduced through QA distractors can lead to noticeable performance degradation in billion-parameter LLMs. TCES with only a single entity modification step combined with adversarial sampling methods and appropriate text embeddings is a feasible way for collecting adversarial entity data. Moreover, TCES ensures that the replacement entity appears natural to humans and is less likely to be detected using grammar check or topical filtering (Morris et al., 2020a). The setting simulates realistic scenarios where the “attacks" may be initiated by unsuspecting users, such as biomedical professionals using an LLM-based clinical decision support system (Liu et al., 2023). In the examples illustrated in Fig. 5 and Appendix E, adversarial distractors can lead to disease misdiagnosis or misprescription of medication, which are detrimental scenarios that may occur in algorithmic decision-making in biomedicine using LLMs. Further improvement on the attack performance may be achieved using learned samplers through rejection mechanisms (Narasimhan et al., 2024) or by finetuning the text embedding to improve the query efficiency of sampling. Our approach may be integrated into interactive platforms for adversarial testing in open-ended human-machine conversation (Chao et al., 2023) for online model auditing.
Future research may focus on how to make LLMs more reliable through prompt engineering (Si et al., 2023; Nori et al., 2023) or leveraging retrieval from nonparametric knowledge bases (Weikum et al., 2021; Soman et al., 2023) or text sources (Wang et al., 2023b) to improve model generalization to long-tail entities. Besides vulnerability identification, the current work motivates adversarial defense strategies based on generalized distances to establish a tiered system for validating adversarial robustness in user-facing and domain-oriented applications.
Broader Impact Statement
The scenarios considered in this work using public datasets relate to the potential real-world deployment of LLMs in biomedical education and clinical decision support. Our results show that the lack of robustness assessment or adversarial testing methods in high-stakes domains may incur serious consequences because of naturally looking adversarial examples. The design and deployment of adversarial testing should be specific to and carried out at every stage of the model development cycle, paired with tools such as adversarially robust model explainability methods for understanding model behavior.
Appendix A Adversarial distractor generation (General case)
@@
$\mathrm{argmax}\,{\Pr}_{\mathrm{LLM}}(Y=\mathcal{K}|\mathcal{Q}(d_{i,j}\rightarrow d_{i,j}^{\prime}))\neq\mathrm{argmax}\,{\Pr}_{\mathrm{LLM}}(Y=\mathcal{K}|\mathcal{Q}).$ (9)
@@ Appendix B Proof of the sampling regimes
Here we present a simple proof of the deterministic limit of the powerscaled distance-weighted sampling described in Eq. 3 of the main text. Although the distance function $h$ is not given a specific form, we assume it to satisfy the basic properties of a metric. Given a perturbation set -$\mathcal{E}_{\textrm{-}k}=\{e_{1},...,e_{m}\}$ with respect to a key $k$, there exists a corresponding set of distance values -$\mathcal{H}_{\textrm{-}k}$.
@@ B.3 The case of $n=0$
When $n=0$, it is straightforward to show that $p_{j}|_{n=0}=1/m$, regardless of $j$ using the definition in Eq. 3. This scenario therefore corresponds to uniform random sampling.
Prompt format Answer the following question without explanation. [[[[Content]]]]: A 23-year-old pregnant woman at 22 weeks gestation presents with burning upon urination. She states it started 1 day ago and has been worsening despite drinking more water and taking cranberry extract. She otherwise feels well and is followed by a doctor for her pregnancy. Her temperature is 97.7°F (36.5°C), blood pressure is 122/77 mmHg, pulse is 80/min, respirations are 19/min, and oxygen saturation is 98% on room air. Physical exam is notable for an absence of costovertebral angle tenderness and a gravid uterus. [[[[Question]]]]: Which of the following are the best treatment for this patient? A: Ampicillin B: Ceftriaxone C: Doxycycline D: Nitrofurantoin [[[[Answer]]]]:
@@ Appendix D Prompt formatting
We structured the prompt with bracketed tags to divide different parts ([Context], [Question], [Answer]). An example taken from the MedQA-USMLE dataset (Jin et al., 2021) is shown below. For the MedMCQA dataset (Pal et al., 2022), only the [Context] and [Answer] [^1] Using the [Context] tag improves model performance than the [Question] tag here. tags were used because the questions are short and generally contain no context. The prompt formatting improves model performance and facilitates answer parsing in the model-generated text. The increased space between paragraphs in the following is only to assist viewing and is not included in the actual prompt.
@@ Appendix E Model evaluation and attack examples
Model inference runtime
The evaluations were carried out on a variety of NVIDIA GPUs depending on the model size. For smaller models such as the Flan-T5 series (Chung et al., 2024), the runtime on each type of question (involving drug or disease mentions) in a dataset was generally between 10 and 30 minutes, while it was from one to more than four hours for larger model such as Flan-UL2 (Tay et al., 2022) and Palmyra-Med-20B (Kamble & Alshikh, 2023). The runtime for each type of questions is between two and four hours.
Figure 6: Quality of blackbox gradient estimation affects the scaling behavior of DiscreteZOO (DZOO) attacks using the CODER and GTE-base text embeddings.
Multi-point gradient estimation
We demonstrate the effects of the quality of gradient estimation on the performance of the blackbox gradient attack (DiscreteZOO (Berger et al., 2021)) using a representative dataset in Fig. 6 using MedAlpaca-7B (Han et al., 2023) and the MedMCQA dataset (Pal et al., 2022). The evaluation used TCES introduced in Algorithm 1. According to Eq. 6, for each entity to attack, we randomly sampled two, five, and ten candidate points in its neighborhood in the embedding space for gradient estimation. The attack on each input instance proceeds with a fixed total of LLM queries (total query budget) as indicated on the horizontal axis. The results show that the larger the number of candidate points used for gradient estimation, the more effective the attack becomes before it consumes the total query budget, as measured by the attack success rate (ASR). Having too few candidate points can result in an early plateau phase in the ASR, which stunts its growth. Yet having more candidate points can slightly reduce the ASR growth in the rapid rise phase of the scaling curve. The general trends are consistent in both the CODER (Yuan et al., 2022) and GTE-base (Li et al., 2023b) embeddings used for the illustration.
Perturbed explaination in adversarial examples
We list some examples of successful entity substitution attacks using adversarial distractors following Fig. 5 in the main text. For each example, the text box on top (the prompt formatting removed) shows the original options on the left side and the perturbed options on the right side. The entity of interest before and after perturbation is colored in red and blue, respectively. The checkmark indicates the model prediction on unperturbed and perturbed questions. Beneath the text box are the corresponding heatmaps of token-wise Shapley values (or Shapley profile) before and after the entity substitution attack. The Shapley profiles for the MedQA-USMLE dataset (Jin et al., 2021) were generated using the model Flan-T5-large and those for the MedMCQA (Pal et al., 2022) using the model Flan-T5-xxl (Chung et al., 2024). The Shapley values were calculated using the partition explainer in shap, which uses the Owen values as an approximation (Chen et al., 2023).
MedQA-USMLE An 18-year-old man presents with a sudden loss of consciousness while playing college football. There was no history of a concussion. Echocardiography shows left ventricular hypertrophy and increased thickness of the interventricular septum. Which is the most likely pathology underlying the present condition? A: Mutation in the myosin heavy chain   ✓ A: Mutation in the myosin heavy chain B: Drug abuse B: Drug abuse C: Viral infection C: Muller barth menger syndrome   ✓ D: Autoimmunity of myocardial fibers D: Autoimmunity of myocardial fibers
@@
MedQA-USMLE A 41-year-old woman comes to the physician because of a 3-month history of anxiety, difficulty falling asleep, heat intolerance, and a 6-kg (13.2-lb) weight loss. The patient’s nephew, who is studying medicine, mentioned that her symptoms might be caused by a condition that is due to somatic activating mutations of the genes for the TSH receptor. Examination shows warm, moist skin and a 2-cm, nontender, subcutaneous mass on the anterior neck. Which of the following additional findings should most raise concern for a different underlying etiology of her symptoms? A: Nonpitting edema   ✓ A: Nonpitting edema B: Atrial fibrillation B: Atrial fibrillation C: Lid lag C: Lid lag D: Fine tremor D: Nonpuerperal galactorrhea   ✓
@@
MedQA-USMLE A 25-year-old man is brought to the emergency department after his girlfriend discovered him at home in a minimally responsive state. He has a history of drinking alcohol excessively and using illicit drugs. On arrival, he does not respond to commands but withdraws all extremities to pain. His pulse is 90/min, respirations are 8/min, and blood pressure is 130/90 mm Hg. Pulse oximetry while receiving bag-valve-mask ventilation shows an oxygen saturation of 95%. Examination shows cool, dry skin, with scattered track marks on his arms and legs. The pupils are pinpoint and react sluggishly to light. His serum blood glucose level is 80 mg/dL. The most appropriate next step in management is intravenous administration of which of the following? A: Naloxone A: Naloxone B: Phentolamine   ✓ B: Phentolamine C: Methadone C: Methadone   ✓ D: Naltrexone D: Tienilic acid
@@
MedQA-USMLE A 70-year-old woman presents to the office for a yearly physical. She states she has recently started experiencing pain in her legs and her back. Last year, she experienced a fracture of her left arm while trying to lift groceries. The patient states that she does not consume any dairy and does not go outside often because of the pain in her legs and back. Of note, she takes carbamazepine for seizures. On exam, her vitals are within normal limits. You suspect the patient might have osteomalacia. Testing for which of the following is the next best step to confirm your suspicion? A: 25-hydroxyvitamin D   ✓ A: 25-hydroxyvitamin D B: 1,25-hydroxyvitamin D B: 1,25-hydroxyvitamin D C: Pre-vitamin D3 C: Pre-vitamin D3   ✓ D: Dietary vitamin D2 D: Stiripentol
@@
MedMCQA Which of the following is granted orphan drug status for the treatment of Dravet’s syndrome? A: Stiripentol A: Stiripentol B: Icatibant B: Tocilizumab   ✓ C: Pre-vitamin D3 C: Pitolisant D: Tafamidis   ✓ D: Tafamidis
@@
MedMCQA Which among the following may be used as a sedatives-hypnotic? A: Zolmitriptan A: Zolmitriptan   ✓ B: Zileuton B: Zileuton C: Zolpidem   ✓ C: N-acetylglucosamine D: Zalcitabine D: Zalcitabine